# FlipedRAG: Black-Box Opinion Manipulation Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 5.4 Mitigation Analysis
In Table 7, we compute the spamicity scores of the adversarial triggers generated by PoisonedRAG and our transfer attack. We then calculate the probability of these adversarial triggers being detected as spam by the spamicity detection system under various thresholds, referred to as the detection rate.
It is evident that, under all threshold conditions, the detection rate of triggers generated by FlipedRAG is significantly lower than those generated by PoisonedRAG. This suggests that directly embedding the question in the document lacks sufficient stealth, making it more easily detectable by simple detection methods. As the detection threshold increases, the detection rate of adversarial triggers gradually decreases, with the most noticeable distinction between the two types of triggers occurring at the 0.2 threshold.
Additionally, our transfer attack utilizes the PAT ranking manipulation strategy, resulting in generated triggers that exhibit key characteristics of PAT, including strong semantic consistency and low text perplexity. These traits enable the triggers to evade detection based on perplexity measures.
TABLE VII: Automatic spamicity detection rates (%) of PoisonedRAG and FlipedRAG.
Threshold 0.3 0.25 0.2 0.15 0.1
PoisonedRAG 29 55 80 93 99
FlipedRAG 4 9 24 51 76
5.5 In-depth Discussion
In this paper, we present a method for transforming the target retrieval model into a white-box model through black-box imitation, thereby facilitating more effective and targeted opinion manipulation of black-box RAG systems. Extensive experiments demonstrate that, in a more realistic black-box setting for RAG systems, attackers can exploit access to the augmented context retrieved by the internal retrieval model. By leveraging this data, attackers can train a surrogate model and create poisoned documents that are both highly effective and stealthy in their ability to launch attacks.
Black-box imitation enhances the effectiveness and stealth of our transfer attack, making it more powerful than common transfer attacks and less likely to be detected. This approach provides a more accurate assessment of the reliability and security of RAG systems. Specifically, the opinion manipulation discussed in this paper focuses on open-ended, controversial topics that can influence human cognition. Consequently, assessing the reliability and security of RAG systems is of significant importance for RAG users, developers, and policymakers.
Our research also has certain limitations. We did not account for the influence of the opinion inherent in the PAT trigger text on RAG responses. However, due to the challenges in controlling the opinion polarity within PAT triggers, we plan to investigate more refined text ranking manipulation techniques in future work.
Given the identified vulnerabilities of RAG models, future research should focus on developing more robust defense strategies. These may include improving the robustness of retrieval algorithms, enhancing the reliability of generation models, and implementing multi-level input filtering mechanisms to mitigate adversarial inputs. Such approaches would enable a balanced optimization of both the understanding and reliability of RAG models.
6 Conclusion
In this paper, we investigate the vulnerability of retrieval-augmented generation (RAG) models to opinion manipulation through black-box attacks on open-ended, controversial topics. We also examine the potential impact of such attacks on user cognition and decision-making. Through systematic experimentation, we introduce a novel adversarial transfer attack strategy that employs black-box imitation to poison the retrieval ranking. This approach significantly alters the polarity of opinions generated by RAG models by crafting adversarial samples, without requiring internal knowledge of the RAG model.
The experimental results demonstrate that our transfer attack strategy effectively alters the opinions expressed in content generated by RAG models, exposing the vulnerability and unreliability of RAG systems when confronted with malicious retrieval corpora. Our proposed strategy is applicable across various dense retrieval models (Contriver, Co-Condenser, and ANCE) as well as different LLMs such as Llama-3, Mixtral, and Vicuna. Additionally, comparative experiments reveal that black-box imitation significantly enhances the effectiveness of transfer attacks in manipulating the opinions embedded in RAG-generated content. Furthermore, the adversarial documents generated using this method exhibit superior stealth compared to other black-box RAG poisoning attacks. To facilitate reproducibility, we have released the codes on the anonymous repository [^1] https://anonymous.4open.science/r/FlipRAG-EB0A/.
More importantly, this form of opinion manipulation could have profound consequences on users’ cognition and decision-making, potentially leading them to accept inaccurate or biased information. This could induce cognitive shifts and distort public opinion, especially on controversial issues, highlighting the critical importance of safeguarding against such vulnerabilities in RAG systems.
7 Ethical Considerations
In conducting research on the vulnerabilities of Retrieval-Augmented Generation (RAG) models, we have meticulously considered and addressed various ethical implications to ensure that our work complies with the highest ethical standards.
Disclosure and Mitigation. A primary ethical concern is the potential for our findings to be exploited by malicious actors in the development of more sophisticated attacks against RAG models. To mitigate this risk, we have implemented several precautionary measures: (1) We have responsibly disclosed the identified vulnerabilities prior to this submission, allowing for timely patching and strengthening of the affected models. Our methodologies, findings, and the ethical considerations involved in this study are thoroughly documented. (2) In addition to identifying vulnerabilities, our research offers comprehensive guidelines and strategies to enhance the security and reliability of RAG models. By providing these solutions, we aim to equip developers and researchers with the tools needed to safeguard against potential misuse.
Data Privacy and User Consent. Our experiments involved utilizing opinion-based experiments and datasets covering a range of topics. To protect the privacy and rights of individuals whose data may be included in these datasets, we have ensured that all datasets are anonymized to eliminate personally identifiable information (PII), thereby safeguarding individual privacy. Furthermore, prior to the commencement of the user experiments, participants were fully informed about the background and purpose of the study, as well as the measures in place to ensure fairness and protect privacy. After the experiment, we reiterated the true data distribution to participants to minimize any potential impact on their cognition. Throughout the research process, we have taken care to conduct all experiments, particularly those involving adversarial attacks, within secure environments to prevent unintended consequences.

### v3

@@ B.5.1. Mitigation Based on Spam Detection
B.5.2. Mitigation Based on Keyword Density
Search Engine Optimization (SEO) is a set of techniques used to improve a website’s visibility and ranking on search engine results pages (SERPs). Keyword density, which is a measure of how often a certain keyword or phrase appears in the content, is a critical factor in SEO as high keyword densities make web pages more relevant to the user query. Keyword stuffing is a straightforward and effective SEO technique that involves the repetition of important keywords on a web page to enhance its search engine ranking. The method employed by PoisonedRAG, which involves adding queries to target documents, constitutes a form of keyword stuffing, as the terms within the queries serve as keywords during the retrieval process.
Numerous SEO detection methods targeting keyword stuffing have been developed, including the aforementioned spam detection based on TF-IDF values. Many scholars have also proposed the use of keyword density as a method for detecting spam and SEO defense. Hui et al. (Hui et al., 2012) suggest that an appropriate keyword density for web pages ranges from 3% to 8%. Consequently, Shahzad(Shahzad et al., 2020) et al. proposed a methodology for identifying spam, which includes a step to check whether the keyword density of a webpage exceeds a certain threshold; if it does, the webpage is excluded as spam. Many search engines and SEO tools classify webpages with keyword density exceeding a certain threshold as spam or compare the keyword density of the target webpage with that of other webpages to determined whether the target webpage is spam. Semrush  [^4] https://www.semrush.com/ compares the webpage’s keyword densities against top-ranking rivals’ keyword densities to see whether there is a disparity.
Many keyword stuffing practices concentrate keywords in specific sections of webpage content (e.g., title tags, anchor text, etc.). Therefore, SEO detection necessitates examining the distribution of keyword density within the content. Tan (Tan et al., 2018) et al. utilized a sliding window approach to derive local word density features in emails, which served as the basis for spam detection. We employed a sliding window technique to analyze keyword density across different sections of the text. If an abnormally high keyword density is detected within any window, it suggests that the document likely contains keyword stuffing.
Given that the RAG attack scenario involves manipulation during the search phase, similar to SEO, we also employed keyword density to detect the RAG attack. Following the method outlined in (Shahzad et al., 2020), we treated the terms in the user query as keywords and calculated the overall keyword density of the document. We calculated the overall keyword density and the maximum keyword density across all sliding windows for documents subjected to the baselines and FlippedRAG attacks, as well as for clean data. The sliding window sizes were set to 20, 50, and 100, with a step size of 5. The experimental results are presented in Table 9.
B.5.3. Mitigation Based on Paraphrasing
Paraphrasing is a widely utilized preprocessing defense originating from computer vision, where a generative model is employed to perform encoding and decoding operations on an adversarial image, thereby mitigating its harmful effects (Jain et al., 2023). In NLP, the paraphrasing defense can alter the text content, thereby changing the adversarial sequence or the attack target, which reduces the success rate of the attack on text models of language models.
Paraphrasing can be applied to both queries and documents. Paraphrasing applied to documents would require processing a large volume of documents, which is impractical in real-world scenarios. Therefore, we implemented paraphrasing on queries. Given a user query $q$, the paraphrasing defense leverages the LLM to rewrite $q$ into $q^{\prime}$, preserving the original semantics while ensuring that $q^{\prime}$ differs as much as possible from $q$. This approach aims to hinder the RAG attack by disrupting the retrieval ranking of poisoned documents associated with specific $q$. We use the following prompt to paraphrase the user query: “Paraphrase the following query and do not to use the words in the original query.”
B.5.4. Mitigation Based on Randomized Mask Smoothing
Randomized smoothing is a defense method aiming at achieving certified robustness against adversarial examples. It constructs a smoothed composite function by introducing random variables to the original function (model), thereby enabling the certification of its robustness under certain conditions. In the field of computer vision(CV), Levine et al. (Levine and Feizi, 2020) proposed a certifiable defense based on randomized smoothing against patch attacks in image classification tasks. This approach involves generating numerous noisy versions of the input and aggregating their classification results to mitigate the impact of adversarial patches, thereby achieving accurate classification. For ranking models, Wu et al. (Wu et al., 2022a) defined the concept of Certified Top-K Robustness for text ranking tasks and developed a smoothed ranking model based on random word substitution to address word substitution ranking attacks, thereby establishing a certifiably robust defense CertDR. CertDR relies on knowledge of the lexicon used by attackers to perform word substitution attacks, whereas Xiang et al. (Xiang et al., 2021) and Zeng et al. (Zeng et al., 2023) explore the use of masking to achieve random smoothing in both CV and NLP domains.
Given that one of the objectives of RAG attacks is to target retrieval ranking models, we explored the use of randomized mask smoothing as a defense mechanism to counteract RAG attacks.
B.5.5. Mitigation Based on RobustRAG
RobustRAG designs keyword-based and decoding-based algorithms for aggregating LLM responses. The framework first processes each retrieved passages constituting the context by combining it with the user query and inputs them individually into the LLM to generate multiple isolated responses. The keyword-based algorithm extracts keywords from these isolated responses and compiles them into a keyword set. Subsequently, the LLM generates the final response to the query based on the extracted keywords with large counts. The decoding-based algorithm aggregates the next-token probability vectors predicted from different isolated passages within the context and generates each subsequent token based on the aggregated probability vector, iteratively producing the final response. Experiments demonstrate that this voting-inspired defense mechanism proves particularly effective against poisoning attacks on factoid questions, consistently limiting the attack success rate to approximately 5% or lower.
For the mitigation approach discussed in this section, we exclusively adopt the keyword-based algorithm, as the decoding-based algorithm exhibits prohibitively long execution times, rendering it currently impractical for real-world deployment.